OpenStudy (anonymous):

Which algebraic property is used to rewrite (3 + m) + y as 3 + (m + y)? Distributive property Commutative property Associative property of multiplication Associative property of addition

OpenStudy (anonymous):

@AG23 @bubbleslove1234 @Carissa15 @dan815 @UsukiDoll @Hero @Xaze @AmyRoseRules

OpenStudy (usukidoll):

well it's not Distributive.

OpenStudy (anonymous):

Yeah i think its the last one

OpenStudy (usukidoll):

it is the last one because the () are placed differently and we have + signs too.

OpenStudy (arindameducationusc):

Yes it is the last one... Good job..!!

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

An operation on 2 numbers a and b is said to be commutative if \[a*b=b*a\] where \[*=+,-,\times,\div\] Clearly addition and multiplication are commutative operations \[a+b=b+a\]\[a \times b = b \times a\] Associative is when \[a*(b*c)=(a*b)*c\] Again, both addition and multiplication are associative